OpenStudy (anonymous):

Factor the trinomial: 3x2 - 7x - 6 (3x + 2)(x - 3) (3x - 2)(x + 3) (3x - 1)(x - 6) (3x + 1)(x - 6)

OpenStudy (anonymous):

please shown steps im really lost:(

OpenStudy (campbell_st):

ok.... here is my method for a quadratic \[ax^2 + bx + c\] multiply a * c in your question its 3 and - 6 3*-6 = -18 find the factors of - 18 that add to -7..... -9, 2 then do this 

OpenStudy (mathstudent55):

A similar method is below: In a polynomial ax^2 + bx + c, find the product ac In your case, 3x^2 - 7x -6, that woud be 3(-6) = -18 Now find two factors of ac that add up to the middle coefficint, b In this case, find two factors of -18 that add up to -7 They are -9 and 2, since -9 * 2 = -18 and -9 + 2 = -7 Nowe break up the middle term, bx, into those two parts In this case it's 3x^2 - 7x - 6 = 3x^2 - 9x + 2x - 6 Now factor by parts: Factor a common factor out of the first two terms, and a common factor out of the last two terms: 3x(x - 3) + 2(x - 3) Now factor out the common factor, (x - 3) (x - 3)(3x + 2)

OpenStudy (anonymous):

Use wolfram alpha